OpenStudy (anonymous):

find the derivative of the function. g(t)=e (to the power of-3/t^2)

zepdrix (zepdrix):

Exponentials of base e are always nice to work with :) Since we know the derivative of e^x = e^x... (d/dx) e^(something messy) = e^(something messy) * (d/dx)(something messy) Understand? :o

OpenStudy (anonymous):

well kinda. i was getting help from this other person and he or she told me to find the derivative of -3/t^2 and then multiply it by e^-3/t^2? right?

OpenStudy (anonymous):

\[\frac{d}{dx}e^{\text{something}}=\frac{d}{dx}\text{something}\times e^{\text{something}}\]

OpenStudy (anonymous):

so this is what i have g prime(t)=0 times t-2 +-2t(-3) is this correct?

OpenStudy (anonymous):

\[\frac{d}{dt}\frac{-3}{t^2}=6t^{-3}=\frac{6}{t^3}\]

OpenStudy (anonymous):

by the power rule right? \[\frac{-3}{t^2}=-3t^{-2}\] so the derivative is \[-2\times -3t^{-2-1}=6t^{-3}=\frac{6}{t^3}\]

OpenStudy (anonymous):

i think i need to start all over again sorry :[

OpenStudy (anonymous):

lol @zepdrix wrote the exact same thing i did, before i did!

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

written in math, the derivative of \(e^{f(t)}\) is \(f'(t)e^{f(t)}\)

zepdrix (zepdrix):

@satellite73 Hey sometimes I see people writing stuff in equation form, but in very large text, how do you do that? :C

OpenStudy (anonymous):

\large

OpenStudy (anonymous):

help me?

OpenStudy (anonymous):

if you ever want to see code, right click and select "show math" as then "latex"

zepdrix (zepdrix):

oh sorry lil XD didn't mean to hijack your post lol

OpenStudy (anonymous):

thanks guys :[

OpenStudy (anonymous):

\[\huge e^x\]

OpenStudy (anonymous):

lotta help ;{{{{{{{{{{

OpenStudy (anonymous):

ok here we have \[e^{f(t)}\] with \(f(t)=-\frac{3}{t^2}, f'(t)=\frac{6}{t^3}\) right?

OpenStudy (anonymous):

or is this still confusing?

OpenStudy (anonymous):

still confusing......im sorry im so HARD to work with

OpenStudy (anonymous):

no problem but we still have the power rule we can use yes?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so our real job is to find the derivative of \(-\frac{3}{t^2}\) right? i mean that is really all we need to do

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

once we have that, we are done

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so how do you write \(-\frac{3}{t^2}\) in exponential notation (in one line, rather than as a fraction?)

OpenStudy (anonymous):

ummm -3(t^-2)

OpenStudy (anonymous):

yes. now apply the all mighty power rule

OpenStudy (anonymous):

okay... so we dont use the product rule at all?

OpenStudy (anonymous):

oh NO now when you have a constant like \(-3\) or any other constant for that matter

OpenStudy (anonymous):

*not

OpenStudy (anonymous):

for example, the derivative of \(\sin(x)\) is \(\cos(x)\) so the derivative of \(7\sin(x)\) is \(7\cos(x)\)

OpenStudy (anonymous):

o okay

OpenStudy (anonymous):

soapplying the power rule to -3(t^-2) its -3(-2t)?

OpenStudy (anonymous):

what is \(-2-1\) ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

yes, so that is your exponent from the power rule subtract one from the exponent of \(-2\) to get \(-2-1=-3\)

OpenStudy (anonymous):

\[\frac{d}{dx}x^n=nx^{n-1}\] whether \(n\) is positive, negative, a fraction, or even an irrational number

OpenStudy (anonymous):

okay so then putting this all together my answer to this question is 6t^-3 *e^-3/t^2? right?

OpenStudy (anonymous):

yes that is the whole story although you might want to write it as \[\frac{6}{t^3}e^{-\frac{3}{t^2}}\]

OpenStudy (anonymous):

okay gotcha thanks SOOOO much for ur help i APPRECIATE IT SO MUCH!

OpenStudy (anonymous):

or not, as the case may be but generally it is helpful not to have negative exponents, especially if you need further computation yw

OpenStudy (anonymous):

invoice is in the mail

OpenStudy (anonymous):

?

OpenStudy (anonymous):

just kidding

OpenStudy (anonymous):

ok :]